OpenStudy (anonymous):

The line y=3x+7 is tangent to the graph y=f(x)at x=4. What is f(4)? What Is f′(4).

OpenStudy (anonymous):

okay let's have some fun.

OpenStudy (anonymous):

f(x) = 7x - x^2 f'(x) = 7- 2x 7- 2x=3 2x= 4 x=2 y = 7x - x^2 y = 14 - 4 y = 10 slope = 3, point (2, 10) y - y1 = m(x - x1) >> slope point equation of the line. y -10 = 3(x - 2) y = 3x - 6+10 y= 3x + 4

OpenStudy (anonymous):

looks like he got it

OpenStudy (tommo_lcfc):

Please explain. Thanks.

OpenStudy (anonymous):

which is f(4) and f'(4)?

OpenStudy (anonymous):

and how did yo uget your first equation? antiderivative?

OpenStudy (anonymous):

nvm

OpenStudy (anonymous):

still an explanation would be nice

OpenStudy (anonymous):

u need it?

OpenStudy (anonymous):

i'll run you through again okay?

OpenStudy (anonymous):

give me a medal :]

OpenStudy (anonymous):

yes plz!

OpenStudy (anonymous):

y=3x+7 is tangent to the graph y=f(x) @ x=4. What's f(4)? What's f'(4)?

OpenStudy (anonymous):

y-f(x)=f'(x)(x-a)

OpenStudy (anonymous):

a=4, f(x)=unknown, f'(x)=unknown

OpenStudy (anonymous):

tangent of graph is found by the formula above

OpenStudy (anonymous):

3x+7=f'(x)(x-a)+f(x)

OpenStudy (anonymous):

tangents have m=-1xrecipricol

OpenStudy (anonymous):

so f(x)=-1/3x+c

OpenStudy (anonymous):

plug in point x=4

OpenStudy (anonymous):

-1/3(4)+c=f(x)

OpenStudy (anonymous):

c being 7?

OpenStudy (anonymous):

or what is c?

OpenStudy (anonymous):

hmm let me do this on paper, i'll refine it

OpenStudy (anonymous):

ahh...you need to integrate

OpenStudy (anonymous):

easy

OpenStudy (anonymous):

dirty tricksters as I thought.... then f'(4) is just 19 right?

OpenStudy (anonymous):

y=19

OpenStudy (anonymous):

it's tricky, i'll show you

OpenStudy (anonymous):

wups did you mean integrate the function you found?

OpenStudy (anonymous):

ahh kk

OpenStudy (anonymous):

@ x=4, y=19

OpenStudy (anonymous):

so you plug that into the integral of tangent

OpenStudy (anonymous):

and find intercept (c)

OpenStudy (anonymous):

c= -33

OpenStudy (anonymous):

how is it -33?

OpenStudy (anonymous):

henceforth, \[f(x)=3/2x^2+7x-33\]

OpenStudy (anonymous):

at the point x=4, y=19 right?

OpenStudy (anonymous):

the tangent is essentially the line of the gradient m or dy/dx

OpenStudy (anonymous):

at that given point (4,19)

OpenStudy (anonymous):

so the line is the integral of the tangent at that point true?

OpenStudy (anonymous):

wait..

OpenStudy (anonymous):

it's messy when you have too many unknowns

OpenStudy (anonymous):

yes i was correct

OpenStudy (anonymous):

okay all the answers: \[f(x)=3/2x^2+7x-33\] f(4)=639 f'(4)=19

OpenStudy (anonymous):

i overcomplicated things a lot sorry